# Functional Eigen-Grasping Using Approach Heatmaps

### v2

@@ V-A Contact and Functional Energy
Contact energy is preferred over force closure metrics for achieving functional grasps. While force closure metrics overly focus on the object’s center of gravity, leading to suboptimal grasps, contact energy offers a more effective method for reliable and functional grasps.
@@ VI-C Real World Evaluation
As illustrated in Fig. 7, the gripper approaches the object based on its pose, grasps it with the non-functional fingers, lifts it, and then presses the functional finger onto the functional part. Our results demonstrate the effectiveness of this approach, with the robotic system successfully lifting all tested objects, as shown in Fig. 8.
@@ VII Conclusion
This paper presents a method for achieving functional grasps across diverse objects using different grippers. By defining the functional components of both the robotic gripper and the object, we integrate a functional constraint to narrow the grasp planning search space, focusing solely on grasps that fulfill the intended function. We introduce the concept of the approach heatmap, which identifies optimal locations on the object for the robotic palm to approach, enabling effective functional grasps. Our experiments demonstrate planning and executing functional grasps on everyday objects using this heatmap.
This method allows each robotic gripper to identify suitable grasping areas, accommodating variations in kinematic design and capabilities. Currently, the approach heatmap generation requires the complete model and specification of the object’s functional part. To generalize this process to novel and previously unseen objects, we aim to automate it using methods like open-vocabulary part segmentation [29], which can detect the functional part from an RGB image. A key challenge will be integrating the concept of functional reachability to generate the heatmap.
Another area for improvement is in integrating the approach heatmap generation and grasp planning processes, as their separation limits the grasp planner’s effectiveness, especially for objects like scissors. Future work will focus on integrating these steps to optimize performance and generalize our method, considering the execution of functional motions. We believe this will enhance the generalization and performance of our method.

### v3

@@ VI-C Real World Evaluation
TABLE II: Grasp Resistance to Force Disturbance and Force Amount to Collapse the Grasp
Object a b c d e f
Weight (g) 361 133 243 1033 148 349
Success / #Trials 30/30 30/30 30/30 30/30 30/30 16/30
Object Deviation (cm) $0.67\pm 0.32$ $0.61\pm 0.38$ $0.89\pm 0.49$ $1.2\pm 0.93$ $0.79\pm 0.43$ $0.52\pm 0.21$
Grasp Collapse Force (N) $6.02\pm 0.51$ $23.54\pm 5$ $20.76\pm 4.1$ $9.69\pm 2.4$ $4.88\pm 0.87$ $4.7\pm 2.3$
@@
As shown in Fig. 7, given the object’s pose, the gripper approaches and secures it with the non-functional fingers, lifts it, and presses the functional finger onto the designated area. Fig. 8 demonstrates the robotic gripper successfully lifting all objects.
We assess the stability of functional grasps, selected based on Table I, using the method from [29]. Specifically, we apply a 4N disturbance force but instead of relying on human observation for scoring, we quantify grasp stability by measuring the average object deviation caused by the disturbance using markers on the object and an RGB camera. Each grasp is tested in over 30 trials per object, with success defined as the gripper maintaining its hold despite applied disturbances. Additionally, we assess the force required to break each grasp, applying up to 30N. These tests are conducted over 10 trials per object, with results summarized in Table II. Among the tested objects, only object F exhibited partial failure, losing its grasp in 14 out of 30 trials.
@@ VII Conclusion
This paper presents a method for achieving functional grasps across diverse objects and grippers by defining functional components of both. We introduce an approach heatmap to identify optimal palm approach locations, narrowing the grasp search space to functionally relevant grasps. Experiments show successful planning and execution on everyday objects using this heatmap.
Currently, heatmap generation relies on the object’s full model and functional specification. To generalize to unseen objects, we aim to automate this with open-vocabulary part segmentation [30]. A key challenge is integrating functional reachability and grasp planning, which limits effectiveness for complex objects like scissors. Future work will unify these processes to enhance generalization and performance.